# Off-policy Evaluation with Deeply-abstracted States

### v2

@@ Appendix B Implementation details
In this section, we present implementation details for forward abstraction (Section B.1) and backward abstraction (Section B.2).
B.1 Implementation details for forward abstraction
We provide details for implementing the proposed forward abstraction in this subsection. We use deep neural networks to parameterize the forward abstraction and estimate the parameters by minimzing the following loss function:
@@ B.1 Implementation details for forward abstraction
By definition, the forward abstraction is required to achieve both model-irrelevance and $\pi$-irrelevance. As discussed in Section 4.2, our approach is to learn a model-irrelevant abstraction, denoted as $\phi$, and then concatenate it with $\{\pi(a|\bullet):a\in\mathcal{A}\}$. We denote the concatenated abstraction by $\phi_{for}$.
@@
where $\mathcal{R}_{\phi_{0}}$ and $\mathcal{T}_{\phi_{0}}$ are the estimated reward and transition functions applied to the abstract state space parameterized by deep neural networks as well, and $|\mathcal{D}|$ is the cardinality of the dataset $\mathcal{D}$.
The inclusion of the third loss function, $\mathcal{L}_{Q}$, is motivated by the demonstrated benefits of utilizing model-free objectives to guide the training of state abstractions in policy learning François-Lavet et al., (2019).
@@
where $Q^{-}$ and $Q$ represent the estimated $Q^{\pi}_{\phi_{for}}$ function applied to the abstract state space during the previous and current iterations, respectively.
The above objectives allow us to effectively train forward abstractions. However, a potential concern is that the resulting abstraction and transition can collapse to some constant $x_{0}$ such that $\phi_{for}(S)\rightarrow x_{0},~{}~{}\forall S\in\mathcal{S}$. To address this limitation, we include the following penalty function of two randomly drawn states to promote diversity in the abstractions:
@@
$\mathcal{L}_{s}=\frac{1}{|\mathcal{D}|}\sum_{(S,S^{\prime})\in\mathcal{D}}C_{1}[\|\phi_{for}(S)-\phi_{for}(S^{\prime})\|_{2}-d_{0}]^{2},$
@@
The forward model architecture is as follow:
@@
B.2 Implementation details for backward abstraction
We provide details for implementing the proposed backward abstraction in this subsection. Similar to Section B.1, we use deep neural networks to parameterize the abstraction $\phi_{back}$ and estimate the parameters by solving the following loss function,
@@ B.2 Implementation details for backward abstraction
Recall that backward-model-irrelevance requires both $\rho^{\pi}$-irrelevance (Definition 6) and (5). To enforce $\rho^{\pi}$-irrelevance, we first introduce behavior-irrelevance, this can be achieved by minimizing the following cross-entropy loss for behavior:
@@
which incorporates history information as mentioned in 4.3. In practice, we do not use all the history information for behaviour policy, instead we use the history up to past two steps: $b(A_{t}|\phi_{1}(S_{t}),A_{t-1},\phi_{1}(S_{t-1}),A_{t-2},\phi_{1}(S_{t-2}))$.
As commented in Section 4.2, (5) holds by satisfying the conditional independence assumption between $(A_{t},\phi(S_{t}))$ and $S_{t+1}$ given $\phi(S_{t+1})$. By Bayesian formula, we can show that it is satisfied by the inverse-model-irrelevance and density-ratio-irrelevance when setting the learning policy $\pi$ to $b$. This motivates us to leverage the two objectives $\mathcal{L}_{inv}$ and $\mathcal{L}_{ratio}$ used by Allen et al., (2021) for training MSA. More details regarding these losses can be found in Section 5 of Allen et al., (2021). Note that to obtain non-sequential states $(s,\tilde{s})$ used in $L_{ratio}$, we flip $s^{\prime}$ in the pairs $(s,s^{\prime})$ in each batch instead of shuffling.
Finally, $\mathcal{L}_{s}$ corresponds to the smoothness penalty introduced in Section B.1. The backward model architecture is:
@@ C.1 Reproducibility
The hyper-parameters to train the proposed forward and backward abstractions can be found in Table B.1.
@@ C.2 Experimental settings and additional results
For the environment LunarLander, we use Adam Kingma and Ba, (2014) optimizer with learning rate $0.003$. Model architectures and hyper-parameters are outlined in B. When conducting OPE, the FQE network has $3$ hidden layers with $64$ nodes per hidden layer for abstraction methods, and is equipped with $5$ hidden layers with $128$ nodes per hidden layer for non-abstracted observations (shown as ‘FQE’ in the plot).
@@
For forward and backward models, we abstract the original state dimension from $300\rightarrow 10$, and for DSA method we reduce dimensions from $300\rightarrow 100\rightarrow 50\rightarrow 20\rightarrow 6$, by [backward, forward, backward, forward] order.
@@
We pre-train an agent by using DQN as our target policy. The agent is trained until there exists an episode that has accumulative discounted rewards exceeding $200$ with discounted rate $\gamma=0.99$. We evaluated oracle value (61.7) of the optimized agent by Monte Carlo method with the same discounted rate. The agent model architecture is as follow:
    DQN(
  (fc1): Linear(in_features=8, out_features=64, bias=True)
  (fc2): Linear(in_features=64, out_features=64, bias=True)
  (fc3): Linear(in_features=64, out_features=4, bias=True)
C.3 Licences for existing assets
We consider the environment from OpenAI Gym (Brockman et al.,, 2016) “LunarLander-v2” with the MIT License and Copyright (c) 2016 OpenAI (https://openai.com).
C.4 Computing resources
To build Figure 5, we trained 3 abstraction methods and one non-abstraction method on 4 different sizes of data, each with 20 runs, under 3 $\epsilon$ values. In average, each run takes approximately 8 minutes for four methods on an E2-series CPU with 64GB memory on GCP. It takes about 32 compute hours to complete all the experiments in the figure.

### v3

@@ Appendix B Implementation details
In this section, we present implementation details for type-I MSA (Section B.1) and type-II MSA (Section B.2).
B.1 Implementation details for type-I MSA
We provide details for implementing the proposed type-I MSA in this subsection. We use deep neural networks to parameterize the type-I MSA and estimate the parameters by minimzing the following loss function:
@@ B.1 Implementation details for type-I MSA
By definition, the type-I MSA is required to achieve both model-irrelevance and $\pi$-irrelevance. As discussed in Section 4.2, our approach is to learn a model-irrelevant abstraction, denoted as $\phi$, and then concatenate it with $\{\pi(a|\bullet):a\in\mathcal{A}\}$. We denote the concatenated abstraction by $\phi_{for}$.
@@
where $\mathcal{R}_{\phi}$ and $\mathcal{T}_{\phi}$ are the estimated reward and transition functions applied to the abstract state space parameterized by deep neural networks, and $|\mathcal{D}|$ is the cardinality of the dataset $\mathcal{D}$.
The inclusion of the third loss function, $\mathcal{L}_{Q}$, is motivated by the demonstrated benefits of utilizing model-free objectives to guide the training of state abstractions in policy learning François-Lavet et al. [2019].
@@
where $Q^{-}$ and $Q$ represent the estimated $Q^{\pi}_{\phi}$ function applied to the abstract state space during the previous and current iterations, respectively.
The above objectives allow us to effectively train type-I MSAs. However, a potential concern is that the resulting abstraction and transition can collapse to some constant $x_{0}$ such that $\phi_{for}(S)\rightarrow x_{0},~{}~{}\forall S\in\mathcal{S}$. To address this limitation, we include the following penalty function of two randomly drawn states to promote diversity in the abstractions:
@@
$\mathcal{L}_{s}=\frac{1}{|\mathcal{D}|}\sum_{(S,S^{\prime})\in\mathcal{D}}C_{1}[\|\phi(S)-\phi(S^{\prime})\|_{2}-d_{0}]^{2},$
@@
The type-I MSA architecture is as follow:
@@
B.2 Implementation details for type-II MSA
We provide details for implementing the proposed type-II MSA in this subsection. Similar to Section B.1, we use deep neural networks to parameterize the abstraction $\phi_{back}$ and estimate the parameters by solving the following loss function,
@@ B.2 Implementation details for type-II MSA
Recall that backward-model-irrelevance requires both behavior-irrelevance (6) and (7). To enforce behavior-irrelevance, we achieve this by minimizing the following cross-entropy loss for behavior:
@@
which incorporates history information (8) as mentioned in Section  4.2. In practice, we do not use all the history information for behaviour policy, instead we use the history up to past two steps: $b(A_{t}|\phi_{1}(S_{t}),A_{t-1},\phi_{1}(S_{t-1}),A_{t-2},\phi_{1}(S_{t-2}))$.
As commented in Section 4.2, (7) holds by satisfying the conditional independence assumption between $(A_{t},\phi(S_{t}))$ and $S_{t+1}$ given $\phi(S_{t+1})$. By Bayesian formula, we can show that it is satisfied by the inverse-model-irrelevance and density-ratio-irrelevance when setting the learning policy $\pi$ to $b$. This motivates us to leverage the two objectives $\mathcal{L}_{inv}$ and $\mathcal{L}_{ratio}$ used by Allen et al. [2021] for training MSA. More details regarding these losses can be found in Section 5 of Allen et al. [2021]. Note that to obtain non-sequential states $(s,\tilde{s})$ used in $L_{ratio}$, we flip $s^{\prime}$ in the pairs $(s,s^{\prime})$ in each batch instead of shuffling.
Finally, $\mathcal{L}_{s}$ corresponds to the smoothness penalty introduced in Section B.1. The type-II MSA architecture is:
@@ C.1 Reproducibility
The hyper-parameters to train the proposed type-I MSA and type-II MSA can be found in Table B.1.
@@ C.2 Experimental settings and additional results
For the environment LunarLander, we use Adam Kingma and Ba [2014] optimizer with learning rate $0.003$, and set the learning rate to be $0.001$ for other environments. Model architectures and hyper-parameters are outlined in Section B. When conducting OPE, the FQE network has $3$ hidden layers with $64$ nodes per hidden layer for abstraction methods, and is equipped with $5$ hidden layers with $128$ nodes per hidden layer for non-abstracted observations (shown as ‘FQE’ in the plot).
C.2.1 Atlantis-ram-v5&SpaceInvaders-ram-v5
Data generating processes
[…]
Model parameters
For type-I and type-II MSAs, we abstract the original state dimension from $128\rightarrow 20$. For Atlantis, DSA method reduces dimensions from $128\rightarrow 80\rightarrow 50\rightarrow 20\rightarrow 8$, by [Type-I, Type-II, Type-I, Type-II] order. For SpaceInvaders, DSA method reduces dimensions from $128\rightarrow 80\rightarrow 50\rightarrow 10\rightarrow 8$, by [Type-I, Type-II, Type-I, Type-II] order.
C.2.2 CartPole-v0
Data generating processes
We manually insert 296 irrelevant features in the state, each following a first order auto-regressive model (AR(1))
$\mathbb{P}(S_{t+1,j}|S_{t},A_{t})=\mathbb{P}(S_{t+1,j}|S_{t,j}),~{}~{}~{}~{}~{}j=5,\dots,300.$
We also define a new state-action-dependent reward as
$\mathcal{R}(s_{t},a_{t})=1-2s_{t,1}^{2}-5s_{t,3}^{2},$
where $s_{t,1}$ and $s_{t,3}$ are the first feature (cart position) and third feature (pole angle) of the state $s_{t}$, to replace the original constant rewards. The number of experiences $n$ in the offline dataset is chosen from $\{200,400,1000,2000\}$. The target policy is determined by the pole angle: we push the cart to the left if the angle is negative and to the right if it is positive. Namely,
$\pi(s_{t})=\mathbb{1}(s_{t,3}>0).$
[…]
Model parameters
For type-I and type-II models, we set the abstracted state dimension as $100$. For DSA method, we apply type-II MSA followed by type-I MSA, reducing the dimension from $300\rightarrow 100\rightarrow 6$.
C.2.3 LunarLander-v2
@@
For type-I and type-II MSAs, we abstract the original state dimension from $300\rightarrow 10$, and for DSA method we reduce dimensions from $300\rightarrow 100\rightarrow 50\rightarrow 20\rightarrow 6$, by [Type-II, Type-I, Type-II, Type-I] order.
@@
We pre-train an agent by using DQN as our target policy. The agent is trained until there exists an episode that has accumulative discounted rewards exceeding $200$ with discounted rate $\gamma=0.99$. We evaluated oracle value (61.7) of the optimized agent by Monte Carlo method with the same discounted rate.
C.3 Computing resources
To build Figure 3, we trained 6 abstraction methods and one non-abstraction method for 4 environments, each on 4 different sizes of data, each with 30 runs, under 3 $\epsilon$ values. In average, each run takes approximately 12 minutes for all models on an 16-core E2-series CPU with 64GB memory on GCP. It takes about 240 compute hours to complete all the experiments in the figure.